Find the volume of the solid by revolving the astroid x⅔+y⅔=a⅔ along x axis

1
Expert's answer
2022-06-01T11:45:00-0400

Due to symmetry, we can consider the region lying in the first quadrant and then multiply the volume of the region by two.

Let's solve the astroid equation for "y^2"


"y^{2\/3}=a^{2\/3}-x^{2\/3}"

"y^2=(a^{2\/3}-x^{2\/3})^3"

"y^2=a^2-3a^{4\/3}x^{2\/3}+3a^{2\/3}x^{4\/3}-x^2"

Hence, the total volume of the solid bounded by the astroid is given by


"V=2\pi\displaystyle\int_{0}^ay^2dx"

"=2\pi\displaystyle\int_{0}^a(a^2-3a^{4/3}x^{2/3}+3a^{2/3}x^{4/3}-x^2)dx"

"=2\pi[a^2x-\dfrac{9}{5}a^{4/3}x^{5/3}+\dfrac{9}{7}a^{2/3}x^{7/3}-\dfrac{1}{3}x^3]\begin{matrix}
 a \\
 0
\end{matrix}"

"=2\pi(a^3-\dfrac{9}{5}a^{4/3}a^{5/3}+\dfrac{9}{7}a^{2/3}a^{7/3}-\dfrac{1}{3}a^3)"

"=\dfrac{32\pi a}{105} ({units}^3)"